OpenStudy (anonymous):

Karen plotted the route of a city tour bus on a coordinate grid. It shows the route as a quadrilateral with vertices at (5, 4), (6, 7), (3, 8), and (2, 5). Which statement about the bus route is true? t is a square of side length 1 unit. It is a square of side length 2sqrt2 units. It is a square of side length sqrt10 units. It is a square of side length 4 units

sam (.sam.):

Side (5,4) >>>>(6,7) = root(10) slope = (4 - 7) / (5 - 6) = 3 Side (6,7)>>>>(3,8) = root(10) slope = (7 - 8) / (6 - 3) = -1/3 Side (3,8)>>>>(2,5 = root(10) slope = (8 - 5) / (3 - 2) = 3 Side (2,5)>>>>(5,4) = root(10) slope = (5 - 4) / (2 - 5) = -1/3

sam (.sam.):

It is a square of side length sqrt10 units.